# Solving hidden monotone variational inequalities with surrogate losses

### v2

@@ Acknowledgments
We thank Nicolas Le Roux, Iosif Sakos, Andrew Patterson, Khurram Javed, and Arushi Jain, for helpful discussions and feedback. This work was supported by Borealis AI through the Borealis AI Global Fellowship Award.
@@ Appendix A Proofs
Let Assumption 3.1 hold, then there exists a fixed stepsize $\eta$ for the the surrogate losses $\{\ell_{t}\}_{t\in\mathbb{N}}$ such that the iterates $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$ produced by Algorithm 1 converge linearly to the solution $z_{\ast}$ if either of the following conditions hold:
$1)\;\;\alpha<\frac{\mu}{L}\,,\qquad\text{or}\qquad 2)\;\;\alpha\leq C\eta^{p}\,,\,C,p>0$
@@
If $\alpha\leq C\eta^{p}$, we require
$1\displaystyle 1-2\eta\mu+2C\eta^{1+p}L+\eta^{2}L^{2}+C\eta^{2p+2}L^{2}<$ 1
[…]
which holds for sufficiently small $\eta$. ∎
@@
where in the inequality we used (17, 20) and in the second inequality $c\geq 2(1+\alpha^{2})$. Taking an expectation given information up to time $t$ and using the expected co-coercivity assumption gives,
@@ Appendix E Applications to Performative Prediction
The prediction dependent shift in targets is represented by the distribution $D(\cdot)$ being a function of the model predictions $g(\theta)$. Note that we have used the formulation of performative prediction with respect to model predictions Mofakhami et al. (2023) instead of parameters.

### v3

@@ Acknowledgments
We thank Nicolas Le Roux, David Kanaa, Iosif Sakos, Andrew Patterson, Khurram Javed, and Arushi Jain, for helpful discussions and feedback. This work was supported by Borealis AI through the Borealis AI Global Fellowship Award.
@@ Appendix A Proofs
Let Assumption 3.1 hold and let $\{z_{t}=g(\theta_{t})\}_{t\in\mathbb{N}}$ be the iterates produced by Algorithm 1. If $\alpha$ and $\eta$ are picked such that $\rho:=1-2\eta(\mu-\alpha L)+(1+\alpha^{2})\eta^{2}L^{2}<1$ then, $z_{t}$ converge linearly to the solution $z_{\ast}$ at the following linear rate:
$\|z_{t+1}-z_{*}\|^{2}\leq\rho^{t}\|z_{1}-z_{*}\|^{2}.$ (15)
Particularly, if $\alpha<\frac{\mu}{L}$ and $\eta<\frac{2(\mu-\alpha L)}{(1+\alpha^{2})L^{2}}$ then $\rho<1$ and if $\alpha\leq\frac{\mu}{2L}$ and $\eta=\frac{2\mu}{5L^{2}}$ then $\rho\leq 1-\frac{\mu^{2}}{5L^{2}}$.
@@
If $\alpha\leq\frac{\mu}{2L}$ and taking $\eta=\frac{2\mu}{5L^{2}}$ we have:
$1-2\eta\mu+2\alpha\eta L+(1+\alpha^{2})\eta^{2}L^{2}$ $\overset{\alpha\leq\nicefrac{{\mu}}{{2L}}\leq\nicefrac{{1}}{{2}}}{\leq}1-\eta\mu+\left(1+\frac{1}{4}\right)\eta^{2}L^{2}$
=$\overset{\eta=\frac{2\mu}{L^{2}}}{=}1-\frac{2\mu^{2}}{5L^{2}}+\frac{\mu^{2}}{5L^{2}}=1-\frac{\mu^{2}}{5L^{2}}.$
@@
where in the inequality we used (19, 22) and in the second inequality $c\geq 2(1+\alpha^{2})$. Taking an expectation given information up to time $t$ and using the expected co-coercivity assumption gives,
@@ Appendix B Non-linear least-squares
Definition B.2 (Polyak (1964); Łojasiewicz (1963)).
A function $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfies the PL condition if there exists $\mu>0$ such that for all $x\in\mathbb{R}^{n}$
$\|\nabla f(x)\|^{2}\geq 2\mu(f(x)-f^{\ast}),$ (23)
where $f^{\ast}=\inf_{x\in\mathbb{R}^{n}}f(x)$.
@@ D.2 The Nonlinear Case
Figure 10: Comparison of average performance of surrogate methods in minimizing the value prediction error for RL tasks with a 16-layer MLP in Ant as measured by outer loop iterations (left) and wallclock time (right). The average value prediction error across 20 runs along with 95% confidence intervals are computed from a fixed test set. The red dashed line represents the lowest value prediction error achieved by any of the algorithms.
D.2.1 Larger network
In Figure 10 we report the average performance of surrogate methods using a 16 layer MLP. We observe similar runtimes and performance to the 2 layer MLP used in Section 4.
@@ Appendix E Applications to Performative Prediction
The prediction dependent shift in targets is represented by the distribution $D(\cdot)$ being a function of the model predictions $g(\theta)$. Note that we have used the formulation of performative prediction with respect to model predictions (Mofakhami et al., 2023) instead of parameters.